OpenStudy (anonymous):

1/sinx+1 +1/cscx+1=1

OpenStudy (dumbcow):

\[\frac{1}{\sin x +1} + \frac{1}{\csc x +1} = 1\] is this correct

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You Keep Asking This Question.

OpenStudy (anonymous):

Nope its different questions.

OpenStudy (anonymous):

csc(x) = 1/sin(x). So, use that and simplify.

OpenStudy (anonymous):

Ohhh Hhahha Dahtt Coo Whaa Yhuu Up Tooo (:

OpenStudy (anonymous):

@Emra- Um I'm getting help on homework..

OpenStudy (anonymous):

Hhahhah Kayyy (:

OpenStudy (anonymous):

@Emra- Yeah this is not a social networking site..lol.

OpenStudy (dumbcow):

\[\frac{1}{\csc x +1} = \frac{1}{\frac{1}{\sin x}+1} = \frac{1}{\frac{1+\sin x}{\sin x}} = \frac{\sin x}{\sin x +1}\]

OpenStudy (anonymous):

Thank you!